OpenStudy (andertj21):

What is the equation of the line that passes through the point (1, 1) and has a y-intercept of 2. x - y + 2 = 0 x + y - 2 = 0 x - y - 2 = 0

OpenStudy (dls):

y=mx+c Passes through (1,1) 1=m+2 m=-1 equation of line=> y=-x+2

OpenStudy (andertj21):

So its B?

OpenStudy (dls):

yes

OpenStudy (andertj21):

What is the equation of a line with a slope of -2 and a y-intercept of 8?